40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 2 19:59:54 2012, Jenne, Update, Alignment, POP, POX, POY, IPPOS, IPANG, REFL all coming out of vac 
    Reply  Fri Nov 2 20:41:53 2012, Jenne, Update, Alignment, AS, REFL camera shots AS_2Nov2012.pngREFL_2Nov2012.png
       Reply  Fri Nov 2 21:40:04 2012, Manasa, Update, Alignment, AS, REFL camera shots 
       Reply  Mon Nov 5 09:53:35 2012, Jenne, Update, Alignment, AS, REFL camera shots AS_5Nov2012.pngREFL_5Nov2012.pngIPPOS_IPANG_weekend_trend.png
    Reply  Sat Nov 3 10:31:59 2012, rana, Update, Alignment, POP, POX, POY, IPPOS, IPANG, REFL all coming out of vac 
Message ID: 7664     Entry time: Fri Nov 2 19:59:54 2012     Reply to this: 7665   7667
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: POP, POX, POY, IPPOS, IPANG, REFL all coming out of vac 

[Evan, Jenne, Jamie]

We used the green laser pointer technique to adjust the POP steering mirrors behind PR2 to get the POP backward beam out onto the table (rather, the mirrors were adjusted so that the green laser pointer, mounted on the POX table, was co-aligned with the beam between PR2 and PR3).

We were unable (why? I feel like it wasn't so hard last time) to see the POX beam, with a camera pointed at an IR card.  We ended up just waving a lens-free CCD camera around on the POX table where we expected POX to be, found the beam, and decided that if the beam was getting to the table, that was good enough.

We then waved the camera around on the POY table, and found the POY beam on the table.  We also moved ITMY up and down in pitch, and saw that the POY beam was moving, so we were satisfied that we had the correct beam.  We should go back and do this same check with POX, although I'm pretty sure that we already have the correct beam.  But checking is good.

We confirmed that IPPOS was coming out of the chambers.  I didn't end up touching any in-vac mirrors for IPPOS, since they all looked centered, and the beam on the table was already centered on the steering mirror on the out-of-vac table.

We got IPANG out of the chamber to the ETMY table.  IPANG has, after the pickoff window, an adjustable mirror, and then a fixed mirror on the BS table.  The beam was very close to the edge, in yaw, on that fixed mirror.  Jamie unclamped it and moved it so the beam was centered, then twisted it until I got beam back down at the end, centered on the first steering mirror down there.  Then Evan and I got the beam centered on the other steering mirror on the in-vac ETMY table, and got the mirror to ~the center of the first out-of-vac steering mirror.  Then Evan adjusted the other steering optics so the beam was hitting the QPD.

We then got the real REFL beam out of the chambers.  I still don't know what that ghost/fake beam is.  Anyhow, we moved PRM around, and saw that the real REFL beam moves, while the fake one doesn't.  We adjusted the adjustable REFL steering mirror in-vac such that the real REFL beam came out to the table.  Once on the AP table, we moved the PRM around again, just to be doubly/triply sure that we had the correct beam.  We put a beam splitter (found on the SP table) after the lens in the REFL path on the AP table, and put the camera on the reflected side of that BS.  This is because, like the AS port, the beam is too dim at the normal camera spot (which for REFL is the transmission through a Y1 mirror).

Jamie has centered IPPOS and IPPANG QPDs, so we should look at the weekend trend come Monday, to see what things look like, and how they drift, if at all.

 

On Monday, we should:

* Check the alignment, and the centering of beams on all mirrors one last time

* Remove all apertures from suspended optics (I think BS and PRM may be the only two that have them at this time)

* Check oplev paths for all mirrors

* Check all pickoffs / beams that need to come out of the vacuum

* Start putting on doors

ELOG V3.1.3-